# A Covariant and Manifestly Projective Invariant Formulation of Thomas-Whitehead Gravity

### v1

@@ 3.1 Projectively Invariant Connection and Schouten Tensor
$\bar{\Gamma}^{a}{}_{bc}$ $=\Gamma^{a}{}_{bc}-\delta^{a}{}_{(b}\text{\textcent}_{c)}=\Pi^{a}_{\ bc}-\delta^{a}_{\ (b}g_{c)}$ (3.1)
where ¢$\text{\textcent}_{c}$ and $g_{c}$ were defined in Eqs. (2.43) and (2.30). Modifying $\Pi^{a}{}_{bc}$ by subtracting the $g_{c}$ term has promoted the connection to affine. At the same time, changing $\Gamma^{a}{}_{bc}$ by subtracting the tensor ¢$\text{\textcent}_{c}$ has not changed its affine character. So $\bar{\Gamma}^{a}{}_{bc}$ indeed transforms as an affine connection, as in Eq. (2.6).
From $\bar{\Gamma}^{a}{}_{bc}$ we can construct a covariant derivative operator ¯$\bar{\nabla}_{a}$, which is manifestly tensorial from the first equality in Eq (3.1) and manifestly projectively invariant from the second. The operation of ¯$\bar{\nabla}_{a}$ is the usual rule for covariant derivatives on a rank-m covariant, rank-n contravariant arbitrary tensor $T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$
¯$\bar{\nabla}_{c}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}=\partial_{c}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ $-\sum_{{\rm p}=1}^{{\rm m}}\bar{\Gamma}^{d}{}_{ca_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}da_{{\rm p}+1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ (3.2)
$+\sum_{{\rm p}=1}^{{\rm n}}\bar{\Gamma}^{b_{{\rm p}}}{}_{cd}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}~{}~{}~{}.$ (3.3)
Thus, substituting the defining Eq. (3.1), we have the following relationship between the three derivatives
¯$\bar{\nabla}_{c}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}=(\nabla_{c}+({\rm m}-{\rm n})\text{\textcent}_{c})T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ $+\sum_{{\rm p}=1}^{{\rm m}}\text{\textcent}_{a_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}ca_{{\rm p}+1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ (3.4)
$-\sum_{{\rm p}=1}^{{\rm n}}\delta^{b_{{\rm p}}}{}_{c}~{}\text{\textcent}_{d}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}$ (3.5)
=(˘$=(\breve{\nabla}_{c}+({\rm m}-{\rm n})g_{c})T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ $+\sum_{{\rm p}=1}^{{\rm m}}g_{a_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}ca_{{\rm p}+1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm n}}}$ (3.6)
$-\sum_{{\rm p}=1}^{{\rm n}}\delta^{b_{{\rm p}}}{}_{c}~{}g_{d}T_{a_{1}\dots a_{{\rm m}}}{}^{b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}.$ (3.7)
With this, it is straightforward to show that the projectively invariant terms listed in (2.83), (2.84), and (2.85) can all be written manifestly so as total derivatives with ¯$\bar{\nabla}_{a}$.
@@
$\bar{\Gamma}^{a}{}_{bc}=$ $\hat{\Gamma}^{a}{}_{bc}+\tilde{C}^{a}{}_{bc}$ (3.8)
$\tilde{C}^{a}{}_{bc}\equiv$ $C^{a}{}_{bc}-\tfrac{1}{{\rm d}+1}\delta^{a}{}_{(b}C^{m}{}_{c)m}~{}~{}~{}.$ (3.9)
This will be an important identity in the coming sections when we derive the field equations as it allows us to use partial integration with ¯$\bar{\nabla}_{a}$ equivalently to ^$\hat{\nabla}_{a}$ their partial integrations differ by a factor of $\tilde{C}^{c}{}_{bc}=0$. This feature forces the traces of the Levi-Civita connection and the projectively invariant connection to be identical. Indeed, tracing Eq. (3.8) and substituting Eq. (2.30) we find that $g_{a}$ is proportional to both the trace of the Levi-Civita connection and the trace of the projectively invariant connection
$g_{a}$ $=-\tfrac{1}{{\rm d}+1}\hat{\Gamma}^{c}{}_{ac}=-\tfrac{1}{{\rm d}+1}\bar{\Gamma}^{c}{}_{ac}~{}~{}~{}.$ (3.10)
We define the tensorial and manifestly projectively invariant tensor $\bar{\mathcal{P}}_{mn}$ by
$\bar{\mathcal{P}}_{mn}=\mathcal{P}_{mn}-\nabla_{m}\text{\textcent}_{n}-\text{\textcent}_{m}\text{\textcent}_{n}=\mathcal{D}_{mn}-\breve{\nabla}_{m}g_{n}-g_{m}g_{n}$ (3.11)
which is tensorial from the first equality and projectively invariant from the second. Comparing this with the first of the terms listed in (2.82) we see it is indeed nothing more than ${\bar{{\mathcal{P}}}}_{cd}$. We construct the trace in the usual manner, namely
$\bar{\mathcal{P}}=$ $g^{mn}\bar{\mathcal{P}}_{mn}=\mathcal{P}-\nabla^{m}\text{\textcent}_{m}-\text{\textcent}_{m}\text{\textcent}^{m}~{}~{}~{}.$ (3.12)
@@
3.2 Projectively Invariant Curvature Tensors
@@
$\bar{R}^{a}_{\ bcd}=$ $\partial_{c}\bar{\Gamma}^{a}{}_{bd}-\partial_{d}\bar{\Gamma}^{a}{}_{bc}+\bar{\Gamma}^{m}{}_{bd}\bar{\Gamma}^{a}{}_{cm}-\bar{\Gamma}^{m}{}_{bc}\bar{\Gamma}^{a}{}_{dm}$ (3.13)
$=$ $R^{a}{}_{bcd}+\delta^{a}{}_{[c}\nabla_{d]}\text{\textcent}_{b}+\delta^{a}{}_{[c}\text{\textcent}_{d]}\text{\textcent}_{b}-\delta^{a}{}_{b}\nabla_{[c}\text{\textcent}_{d]}$ (3.14)
$=$ $\mathcal{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\breve{\nabla}_{d]}g_{b}+\delta^{a}{}_{[c}g_{d]}g_{b}$ (3.15)
which is manifestly covariant and projectively invariant when expressed as in the first line. In contrast, it is only manifestly covariant when expressed as in the second line and only manifestly projectively invariant when expressed as in the third line. We will express the other curvature tensors in this section in the same way: manifestly covariant and projectively invariant in their first lines, manifestly only one or the other in each of their other lines. The derivative ˘$\breve{\nabla}_{d}g_{b}$ is defined as in Eq. (2.16). It is worth noting that all three of the curvature tensors are commutators of derivative operators, and so are antisymmetric in their last two indices. However, the lack of metric compatibility means they are not generally antisymmetric in the first two indices (i.e. $R_{abcd}\neq-R_{bacd}$). Proceeding in the usual manner, we also construct the associated Ricci tensors by contracting over the first and third indices
@@
$=$ $\mathcal{R}_{bd}+({\rm d}-1)\breve{\nabla}_{b}g_{d}+({\rm d}-1)g_{b}g_{d}~{}~{}~{}.$ (3.16)
@@
$=$ $\mathcal{R}+({\rm d}-1)\breve{\nabla}^{b}g_{b}+({\rm d}-1)g^{b}g_{b}$ (3.17)
where raising/lowering is defined as occurring outside of the derivative $\nabla^{b}\equiv g^{bd}\nabla_{d}$ and ˘$\breve{\nabla}^{b}\equiv g^{bd}\breve{\nabla}_{d}$ and $g^{b}\equiv g^{bd}g_{d}$.
@@
$=$ $\mathcal{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\mathcal{D}_{d]b}~{}~{}~{}.$ (3.18)
The projective Ricci tensor is built by contracting $\mathcal{K}^{a}_{\ bcd}$ over the first and third indices:
$\mathcal{K}_{mn}=$ $\bar{R}_{mn}+(d-1)\bar{\mathcal{P}}_{mn}$
$=$ $R_{mn}+(d-1)\mathcal{P}_{mn}-d\mathcal{P}_{[mn]}$
$=$ $\mathcal{R}_{mn}+(d-1)\mathcal{D}_{mn}$ (3.19)
which is necessarily symmetric. We also form the projective Ricci scalar by taking the trace as
$\mathcal{K}=g^{mn}\mathcal{K}_{mn}=$ $\bar{R}+(d-1)\bar{\mathcal{P}}$
$=$ $R+(d-1)\mathcal{P}$
$=$ $\mathcal{R}+(d-1)\mathcal{D}~{}~{}~{}.$ (3.20)
@@
$=$ ˘$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}~{}~{}~{}.$ (3.21)
In summary, we see each of the above curvature tensors are manifestly covariant and tensor in their first lines, where they are each written entirely in terms of the barred objects.
3.3 Covariant and Manifestly Projective Invariant TW Field Equations

### v3

@@ III.1 Projectively Invariant Connection and Schouten Tensor
$\bar{\Gamma}^{a}{}_{bc}$ $=\Gamma^{a}{}_{bc}-\delta^{a}{}_{(b}\text{\textcent}_{c)}=\Pi^{a}_{\ bc}-\delta^{a}_{\ (b}g_{c)}$ (104)
where ¢$\text{\textcent}_{c}$ and $g_{c}$ were defined in Eqs. (49) and (36). Modifying $\Pi^{a}{}_{bc}$ by subtracting the $g_{c}$ term has promoted the connection to affine. At the same time, changing $\Gamma^{a}{}_{bc}$ by subtracting the tensor ¢$\text{\textcent}_{c}$ has not changed its affine character. So $\bar{\Gamma}^{a}{}_{bc}$ indeed transforms as an affine connection, as in Eq. (6).
From $\bar{\Gamma}^{a}{}_{bc}$ we can construct a covariant derivative operator ¯$\bar{\nabla}_{a}$, which is covariant from the first equality in Eq (104) and manifestly projective invariant from the second. The operation of ¯$\bar{\nabla}_{a}$ is the usual rule for covariant derivatives on a rank-m covariant, rank-n contravariant arbitrary tensor $T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}$
¯$\bar{\nabla}_{c}T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}=$ $\partial_{c}T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}-\sum_{{\rm p}=1}^{{\rm m}}\bar{\Gamma}^{d}_{\ ca_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}da_{{\rm p}+1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}+\sum_{{\rm p}=1}^{{\rm n}}\bar{\Gamma}^{b_{{\rm p}}}_{\ cd}T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}~{}.$ (105)
Thus, substituting the defining Eq. (104), we have the following relationship between the three derivatives
¯$\bar{\nabla}_{c}T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}=$ $(\nabla_{c}+({\rm m}-{\rm n})\text{\textcent}_{c})T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}+\sum_{{\rm p}=1}^{{\rm m}}\text{\textcent}_{a_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}ca_{{\rm p}+1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}-\sum_{{\rm p}=1}^{{\rm n}}\delta^{b_{{\rm p}}}_{\ \ c}~{}\text{\textcent}_{d}T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}$ (106)
$=$ (˘$(\breve{\nabla}_{c}+({\rm m}-{\rm n})g_{c})T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}+\sum_{{\rm p}=1}^{{\rm m}}g_{a_{{\rm p}}}T_{a_{1}\dots a_{{\rm p}-1}ca_{{\rm p}+1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ b_{1}\dots b_{{\rm n}}}-\sum_{{\rm p}=1}^{{\rm n}}\delta^{b_{{\rm p}}}_{\ \ c}~{}g_{d}T_{a_{1}\dots a_{{\rm m}}}^{\ \ \ \ \ \ b_{1}\dots b_{{\rm p}-1}db_{{\rm p}+1}\dots b_{{\rm n}}}.$ (107)
With this, it is straightforward to show that the projectively invariant terms listed in (99), (101), and (103) can all be written manifestly so as total derivatives with ¯$\bar{\nabla}_{a}$.
@@
$\bar{\Gamma}^{a}_{\ bc}=$ $\hat{\Gamma}^{a}_{\ bc}+\tilde{C}^{a}_{\ bc}$ (108)
$\tilde{C}^{a}_{\ bc}\equiv$ $C^{a}_{\ bc}-\tfrac{1}{{\rm d}+1}\delta^{a}_{\ (b}C^{m}_{\ \ c)m}$ (109)
$=$ $C^{a}{}_{bc}-\delta^{a}_{\ (b}\text{\textcent}_{c)}~{}.$ (110)
This will be an important identity in the coming sections when we derive the field equations as it allows us to use partial integration with ¯$\bar{\nabla}_{a}$ equivalently to ^$\hat{\nabla}_{a}$ their partial integrations differ by a factor of $\tilde{C}^{c}{}_{bc}=0$. This feature forces the traces of the Levi-Civita connection and the projectively invariant connection to be identical. Indeed, tracing Eq. (108) and substituting Eq. (36) we find that $g_{a}$ is proportional to both the trace of the Levi-Civita connection and the trace of the projectively invariant connection
$g_{a}$ $=-\tfrac{1}{{\rm d}+1}\hat{\Gamma}^{c}{}_{ac}=-\tfrac{1}{{\rm d}+1}\bar{\Gamma}^{c}{}_{ac}~{}.$ (111)
We define the tensorial and manifestly projective invariant tensor $\bar{\mathcal{P}}_{mn}$ by
$\bar{\mathcal{P}}_{mn}=$ $\mathcal{P}_{mn}-\nabla_{m}\text{\textcent}_{n}-\text{\textcent}_{m}\text{\textcent}_{n}$ (112)
$=$ $\mathcal{D}_{mn}-\breve{\nabla}_{m}g_{n}-g_{m}g_{n}$ (113)
which is tensorial from the first equality and projectively invariant from the second. Comparing this with the first of the terms listed in (97) we see it is indeed nothing more than ${\bar{{\mathcal{P}}}}_{cd}$. We construct the trace in the usual manner, namely
$\bar{\mathcal{P}}=$ $g^{mn}\bar{\mathcal{P}}_{mn}=\mathcal{P}-\nabla^{m}\text{\textcent}_{m}-\text{\textcent}_{m}\text{\textcent}^{m}~{}.$ (114)
@@
III.2 Projectively Invariant Curvature Tensors
@@
$\bar{R}^{a}_{\ bcd}=$ $\partial_{c}\bar{\Gamma}^{a}{}_{bd}-\partial_{d}\bar{\Gamma}^{a}{}_{bc}+\bar{\Gamma}^{m}{}_{bd}\bar{\Gamma}^{a}{}_{cm}-\bar{\Gamma}^{m}{}_{bc}\bar{\Gamma}^{a}{}_{dm}$ (115)
$=$ $R^{a}{}_{bcd}+\delta^{a}{}_{[c}\nabla_{d]}\text{\textcent}_{b}+\delta^{a}{}_{[c}\text{\textcent}_{d]}\text{\textcent}_{b}-\delta^{a}{}_{b}\nabla_{[c}\text{\textcent}_{d]}$ (116)
$=$ $\mathcal{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\breve{\nabla}_{d]}g_{b}+\delta^{a}{}_{[c}g_{d]}g_{b}$ (117)
which is covariant and manifestly projectively invariant when expressed as in the first line. In contrast, it is only covariant in each of its terms when expressed as in the second line and only manifestly projective invariant when expressed as in the third line. We will express the other curvature tensors in this section in the same way: covariant and manifestly projectively invariant in their first lines, manifestly only one or the other in each of their other lines. The derivative ˘$\breve{\nabla}_{d}g_{b}$ is defined as in Eq. (19). It is worth noting that all three of the curvature tensors are commutators of derivative operators, and so are antisymmetric in their last two indices. However, the lack of metric compatibility means they are not generally antisymmetric in the first two indices (i.e. $R_{abcd}\neq-R_{bacd}$). Proceeding in the usual manner, we also construct the associated Ricci tensors by contracting over the first and third indices
@@
$=$ $\mathcal{R}_{bd}+({\rm d}-1)\breve{\nabla}_{b}g_{d}+({\rm d}-1)g_{b}g_{d}~{}.$ (118)
@@
$=$ $\mathcal{R}+({\rm d}-1)\breve{\nabla}^{b}g_{b}+({\rm d}-1)g^{b}g_{b}$ (119)
where raising/lowering is defined as occurring outside of the derivative and the raised $g^{b}$ is defined as follows
$\nabla^{b}\equiv g^{bd}\nabla_{d}~{}~{}~{},~{}~{}~{}\breve{\nabla}^{b}\equiv g^{bd}\breve{\nabla}_{d}~{}~{}~{},~{}~{}~{}g^{b}\equiv g^{bd}g_{d}~{}.$ (120)
@@
$=$ $\mathcal{R}^{a}{}_{bcd}+\delta^{a}{}_{[c}\mathcal{D}_{d]b}~{}.$ (121)
Contraction over the first and third indices as before yields the projective Ricci tensor. In terms of the three curvatures this is
$\mathcal{K}_{mn}=$ $\bar{R}_{mn}+({\rm d}-1)\bar{\mathcal{P}}_{mn}$
$=$ $R_{mn}+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}\mathcal{P}_{[mn]}$
$=$ $\mathcal{R}_{mn}+({\rm d}-1)\mathcal{D}_{mn}$ (122)
which is necessarily symmetric. Contraction with the metric again yields the projective Ricci scalar. In terms of the three curvatures this is
$\mathcal{K}=$ $\bar{R}+({\rm d}-1)\bar{\mathcal{P}}$
$=$ $R+({\rm d}-1)\mathcal{P}$
$=$ $\mathcal{R}+({\rm d}-1)\mathcal{D}~{}.$ (123)
@@
$=$ ˘$\breve{\nabla}_{[c}\mathcal{D}_{d]b}+g_{a}\mathcal{K}^{a}{}_{bcd}~{}.$ (124)
In summary, we see each of the above curvature tensors are covariant and manifestly projective invariant in their first lines, where they are each written entirely in terms of the barred objects.
It will become useful to expand these tensors in terms of the Levi-Civita connection as
$\bar{R}^{a}_{\ bcd}=$ $\hat{R}^{a}_{\ bcd}+\tilde{Q}^{a}_{\ bcd}$ (125)
$\bar{R}_{mn}=$ $\hat{R}_{mn}+\tilde{Q}_{mn}$ (126)
$\bar{R}=$ $\hat{R}+\tilde{Q}$ (127)
where we have introduced
$\tilde{Q}^{a}_{\ bcd}\equiv$${\hat{\nabla}}_{[c}{\tilde{C}}^{a}_{\ d]b}-{\tilde{C}}^{e}_{\ b[c}{\tilde{C}}^{a}_{\ d]e}$ (128)
$\tilde{Q}_{mn}\equiv\tilde{Q}^{a}_{\ man}=$${\hat{\nabla}}_{a}{\tilde{C}}^{a}_{\ mn}-{\tilde{C}}^{b}_{\ ma}{\tilde{C}}^{a}_{\ nb}$ (129)
$\tilde{Q}\equiv g^{mn}\tilde{Q}_{mn}=$${\hat{\nabla}}_{a}{\tilde{C}}^{am}_{\ \ \ m}-{\tilde{C}}^{b}_{\ ma}{\tilde{C}}^{am}_{\ \ \ b}~{}.$ (130)
In terms of ¯${\bar{\nabla}}_{a}$, these take the following form
$\tilde{Q}^{a}_{\ bcd}=$${\bar{\nabla}}_{[c}{\tilde{C}}^{a}_{\ d]b}+{\tilde{C}}^{e}_{\ b[c}{\tilde{C}}^{a}_{\ d]e}$ (131)
$\tilde{Q}_{mn}=$${\bar{\nabla}}_{a}{\tilde{C}}^{a}_{\ mn}+{\tilde{C}}^{b}_{\ ma}{\tilde{C}}^{a}_{\ nb}$ (132)
$\tilde{Q}=$${\bar{\nabla}}_{a}{\tilde{C}}^{am}_{\ \ \ m}-{\tilde{C}}^{b}_{\ ma}{\tilde{C}}^{am}_{\ \ \ b}~{}.$ (133)
Note that $\tilde{Q}$ can be written with ^${\hat{\nabla}}_{a}$ and ¯${\bar{\nabla}}_{a}$ interchangeable with no other modification and no change in signs. This is because the divergence of an arbitrary vector is the same in terms of either ^${\hat{\nabla}}_{a}$ and ¯${\bar{\nabla}}_{a}$ as we will prove in the next section. This feature will allow us to integrate by parts Lagrangians with either ^${\hat{\nabla}}_{a}$ or ¯${\bar{\nabla}}_{a}$ as we will see shortly.
In terms of $\tilde{Q}^{a}_{\ bcd}$ and $Q^{a}_{\ bcd}$ and their associated Ricci tensors and Ricci scalars, Eqs. (III.2), (III.2), and (III.2) take the form
$\mathcal{K}^{a}{}_{bcd}=$ $\hat{R}^{a}_{\ bcd}+\tilde{Q}^{a}_{\ bcd}+\delta^{a}{}_{[c}\bar{\mathcal{P}}_{d]b}$ (134)
$=$ $\hat{R}^{a}_{\ bcd}+Q^{a}_{\ bcd}+\delta^{a}{}_{[c}\mathcal{P}_{d]b}-\delta^{a}{}_{b}\mathcal{P}_{[cd]}$ (135)
$\mathcal{K}_{mn}=$ $\hat{R}_{mn}+\tilde{Q}_{mn}+({\rm d}-1)\bar{\mathcal{P}}_{mn}$ (136)
$=$ $\hat{R}_{mn}+Q_{mn}+({\rm d}-1)\mathcal{P}_{mn}-{\rm d}\mathcal{P}_{[mn]}$ (137)
$\mathcal{K}=$ $\hat{R}+\tilde{Q}+({\rm d}-1)\bar{\mathcal{P}}$ (138)
$=$ $\hat{R}+Q+({\rm d}-1)\mathcal{P}~{}~{}~{}.$ (139)
These relationships will be useful when we compare the forms of the equations of motion across different formulations.
III.3 Covariant and Manifestly Projective Invariant TW Field Equations